# On DESI’s DR2 exclusion of CDM

### v1

@@ Abstract
The DESI collaboration, combining their Baryon Acoustic Oscillation (BAO) data with cosmic microwave background (CMB) anisotropy and supernovae data, have found significant indication against CDM cosmology. The significance of the exclusion of CDM can be interpreted entirely as significance of the detection of the parameter that measures variation of the dark energy equation of state. We emphasize that DESI’s DR2 exclusion of CDM is quoted in the articles for a combination of BAO and CMB data with each of three different and overlapping supernovae datasets (at 2.8-sigma for Pantheon+, 3.8-sigma for Union3, and 4.2-sigma for DESY5). We show that one can neither choose amongst nor average over these three different significances. We demonstrate how a principled statistical combination yields a combined exclusion significance of 3.1-sigma. Further we argue that, based on available knowledge, and faced with these competing significances, the most secure inference from the DESI DR2 results is the 3.1-sigma level exclusion of CDM, obtained from combining DESI+CMB alone, while omitting supernovae.
@@
The DESI collaboration has additionally carried out further more detailed investigation of dark energy implications (Calderon et al., 2024; Lodha et al., 2025a, b; Gan et al., 2025), primarily in support of the key paper result. There is already a massive literature following the very similar DESI DR1 result (Adame et al., 2025). This literature contains a wide range of views, both supportive and critical, most of which continue to apply to DR2. The interpretation of the results, therefore, remains very much an open issue.
All the collaborations discussed here have put stupendous efforts into data acquisition and analysis, providing the cosmology community with remarkable resources. We take these entire processes, leading up to the construction of data likelihood functions, at face value, and address only issues of the final interpretation of the results obtained.
@@ 2 The nature of the exclusion
There is the inconvenience that the pivot scale depends on the choice of datasets, though in practice the dependence is not strong (Adame et al., 2025; Abdul-Karim et al., 2025). It would for instance be a step forward for everyone to agree a common pivot scale, in the same way the large-scale structure community choose to quote , the perturbations on scale . Pivoting at (), for instance, would be quite suitable for minimizing the unnecessary loss of accuracy when presenting marginalized results – see for instance Francis et al. (2007).
@@ 3.1 Combining exclusion significance over methodological differences
Now imagine the opposite regime, where the same input data are analysed by three teams under different methodologies, giving consistent but somewhat different results. It’s the same data, so we cannot combine by multiplying. All we can do is choose a probability for each analysis being correct, for example with equal weights. Now the parameter probability distributions add, not multiply
@@ 3.2 Implications for DESI DR2
It is tempting to instead average the three quoted sigmas, or take the middle one as indicative. But this is not what principled statistical inference implies. Coincidentally (or perhaps not entirely so), 3.1-sigma is also the quoted exclusion for DESI+CMB without any supernova data!
@@ 4 The supernovae compilations
Having reached our conclusion on overall exclusion significance, we now study the samples for trends and obstacles to future combination. Let’s take a closer look.
@@ 4.1 DESY5, high- and low-redshift
Prompted by Efstathiou (2024)’s uncovering of a possible calibration mismatch in the low-redshift common DESY5/Pantheon+ sample (see also Huang et al. (2025) and our next subsection), Notari et al. (2024) constructed a ‘split’ DESY5 likelihood that allows the relative calibration of the low- and high-redshift subsamples to float. This essentially removes the anchoring effect of the low-redshift sample and weakens the exclusion (to 1.7-sigma for an analysis using the DESI DR1 BAO data, after marginalising over the relative offset). This mimics the effect of removing the low-redshift sample entirely.
@@ 4.2 Pantheon+ versus DESY5
However, the ultimate result is quite different depending on whether the overlapping SNe are excised from the DESY5 sample or from Pantheon+ (Notari et al., 2024). If the reduced DESY5 dataset is added to Pantheon+, the exclusion remains at 2.5-sigma as for Pantheon+ alone (remember this is for DESI DR1). If instead the reduced Pantheon+ dataset is added to DESY5, the exclusion of 3.8-sigma almost matches the 3.9-sigma from DESY5 alone. In each case the addition of the extra data does improve the overall constraining power, but the best-fit also shifts in such a way that the exclusion of CDM is practically unchanged.
@@ 4.3 Pantheon+ versus Union3
In the Union3 analysis of Rubin et al. (2023), they note that they end up with higher uncertainties than Pantheon+, despite having a larger sample which contains nearly all of Pantheon+. They point to specific additional uncertainties they include which are responsible for this, most importantly intergalactic extinction and the redshift evolution of correlations with host-galaxy properties. This creates a case that the Pantheon+ uncertainties may have some systematic under-estimation.222Separately, Kim (2024) has noted some inconsistency of priors in DESI’s use of the Union3 likelihood in the DR1 release, but did not find significant impact.
@@ Acknowledgements
We thank Ed Copeland and Martin White for discussions. This work was supported by the Fundação para a Ciência e a Tecnologia (FCT) through the research grants UIDB/04434/2020 and UIDP/04434/2020. M.C. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02581/2018 and POPH/FSE (EC). A.R.L. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02854/2017 and POPH/FSE (EC). This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology). This work benefited from the EuCAPT rapid response workshop on April 14th 2025; particular thanks to Kyle Dawson, Colin Hill, and Eiichiro Komatsu.

### v3

@@ Abstract
The DESI collaboration, combining their Baryon Acoustic Oscillation (BAO) data with cosmic microwave background (CMB) anisotropy and supernovae data, have found significant indication against the CDM cosmology. This can also be interpreted as the significance of the detection of the parameter that measures variation of the dark energy equation of state. DESI’s DR2 article quotes exclusion of CDM for combinations of BAO and CMB data with each of three different and overlapping supernovae compilations (at 2.8-sigma for Pantheon+, 3.8-sigma for Union3, and 4.2-sigma for DESY5). We show that one can neither choose amongst nor average over these three different significances. We demonstrate how a principled statistical combination yields a combined exclusion significance of 3.1-sigma. Further we argue that, faced with these competing significances, the most secure inference from the DESI DR2 results is the 3.1-sigma level exclusion of CDM obtained from combining DESI+CMB alone, omitting supernovae.
@@
The DESI collaboration have additionally carried out further more detailed investigation of dark energy implications (Calderon et al., 2024; Lodha et al., 2025a, b; Gan et al., 2025), primarily in support of the key paper result. There is already a massive literature following the very similar DESI DR1 result (Adame et al., 2025). This literature contains a wide range of views, both supportive and critical, most of which continue to apply to DR2. The interpretation of the results, therefore, remains very much an open issue.
All the collaborations discussed here have put outstanding efforts into data acquisition and analysis, providing the cosmology community with remarkable resources. We take these entire processes, leading up to the construction of data likelihood functions, at face value, and address only issues of the final interpretation of the results obtained.
@@ 2 The nature of the exclusion
There is the inconvenience that the pivot scale is dependent on the choice of datasets, though in practice the dependence is not strong (Adame et al., 2025; Abdul-Karim et al., 2025). It would for instance be a step forward for teams and collaborations to agree a common pivot scale, in the same way the large-scale structure community choose to quote , the perturbations on scale . Pivoting at (), for instance, would be quite suitable for minimizing the unnecessary loss of accuracy when presenting marginalized results – see for instance Francis et al. (2007).
@@ 3.1 Combining exclusion significance over methodological differences
Now imagine the opposite regime, where the same input data are analysed by three teams under different methodologies, giving consistent but somewhat different results. It’s the same data, so we cannot combine by multiplying. All we can do is choose a probability for each analysis being correct, for example with equal weights. Now the parameter posterior probability distributions add, not multiply
@@
The weighted combination of the posteriors seen in equation (2) bears some resemblence to Bayesian Model Averaging (see Parkinson & Liddle (2013) for a review), but there the dataset is fixed and different models featuring common parameters are combined to give a single posterior on a shared parameter. In that case posterior model weights are used in the averaging, but that is not possible here as there is no way to assign posterior probabilities to each dataset choice. Another related technique is hyperparameters (Hobson et al., 2002), which can be used to set a probability that a given dataset within a compilation of independent data is correct given specific model assumptions, but that again is not available here because the datasets in question are not independent.
@@ 3.2 Implications for DESI DR2
It is tempting to instead average the three quoted sigmas, or take the middle one as indicative. But this is not what principled statistical inference implies. Coincidentally (or perhaps not entirely so), 3.1-sigma is also the quoted exclusion for DESI+CMB without any supernova data.222Adding the supernovae does significantly tighten the overall constraint in the – plane, principally by cutting off the regions furthest from CDM. It is the exclusion level of CDM that does not change.
@@ 4 The supernovae compilations
Having reached our conclusion on overall exclusion significance, we now study the samples for trends and obstacles to future combination. It’s time for a closer look.
@@ 4.1 DESY5, high- and low-redshift
Prompted by Efstathiou (2025)’s uncovering of a possible calibration mismatch in the low-redshift common DESY5/Pantheon+ sample (see also Huang et al. (2025) and our next subsection), Notari et al. (2025) constructed a ‘split’ DESY5 likelihood that allows the relative calibration of the low- and high-redshift subsamples to float. This essentially removes the anchoring effect of the low-redshift sample and weakens the exclusion (to 1.7-sigma for an analysis using the DESI DR1 BAO data, after marginalising over the relative offset). We wish to empasise here that this mimics removing the low-redshift sample entirely.
@@ 4.2 Pantheon+ versus DESY5
However, the ultimate result is quite different depending on whether the overlapping SNe are excised from the DESY5 sample or from Pantheon+ (Notari et al., 2025). If the reduced DESY5 dataset is added to Pantheon+, the exclusion remains at 2.5-sigma as for Pantheon+ alone (remember this is for DESI DR1). If instead the reduced Pantheon+ dataset is added to DESY5, the exclusion of 3.8-sigma almost matches the 3.9-sigma from DESY5 alone. In each case the addition of the extra data does improve the overall constraining power, but the best-fit also shifts in such a way that the exclusion of CDM is practically unchanged.
@@ 4.3 Pantheon+ versus Union3
In the Union3 analysis of Rubin et al. (2023), they note that they end up with higher uncertainties than Pantheon+, despite having a larger sample which contains nearly all of Pantheon+. They point to specific additional uncertainties they include which are responsible for this, most importantly intergalactic extinction and the redshift evolution of correlations with host-galaxy properties. This creates a case that the Pantheon+ uncertainties may have some systematic under-estimation.333Separately, Kim (2024) has noted a difference between DESI’s priors and those used to derive the Union3 spline-interpolated posterior distribution in the DESI DR1 release, but did not find significant impact.
@@ Acknowledgements
We thank Ed Copeland, George Efstathiou, Alex Kim, and Martin White for discussions. This work was supported by the Fundação para a Ciência e a Tecnologia (FCT) through the research grants UIDB/04434/2020 and UIDP/04434/2020. M.C. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02581/2018 and POPH/FSE (EC). A.R.L. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02854/2017 and POPH/FSE (EC). This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology). This work benefited from the EuCAPT rapid response workshop on April 14th 2025; particular thanks to Kyle Dawson, Colin Hill, and Eiichiro Komatsu.